OpenStudy (anonymous):

Name the y value of the solution to the system of equations. Use substitution: -4x + y = 6 -5x - y = 21

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

please help

OpenStudy (owlcoffee):

all you do is take one of the equations, solve for one variable, and replace it in the other, I'll take the first one and solve for "x": \[-4x+y=6\] \[x=-\frac{ 6-y }{ 4 }\] And replace the value in the second equation: \[-5x-y=21\] \[-5(-\frac{ 6-y }{ 4 })-y=21\] and then simpify it a little: \[\frac{ 5(6-y)-4y }{ 4 }=21\] \[5(y-6)-4y=84\] all you have to do from here on, is solve for y.

OpenStudy (anonymous):

i still need help lol how do you solve for y

OpenStudy (owlcoffee):

When you solve for y, you have to isolate it on the left or right side of the equal sign: \[5(y-6)-4y=84\] \[5y-30-4y=84\] so, operating the terms with the same variable y and simplifying: \[y-30=84\] \[y=114\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

is x not (y-6)/4 ??

OpenStudy (anonymous):

i need help with more problems

OpenStudy (anonymous):

Determine the number of solutions for the system of equations. 2x + 8y = 6 -5x - 20y = -15

OpenStudy (anonymous):

sorry did not notice the minus

OpenStudy (anonymous):

Determine the number of solutions for the system of equations. 2x + 8y = 6 -5x - 20y = -15 A. No Solutions B. One Solution C. Exactly 2 solutions D. Infinite number of solutions

OpenStudy (anonymous):

Josh and his cousin Lucy are picking apples in their grandparents' orchard. Josh has filled 13 baskets with apples and is filling them at a rate of 3 baskets per hour. Lucy has 9 full baskets and will continue picking at 5 baskets per hour. Once the cousins get to the point where they have filled the same number of baskets, they will carry them to the barn and then eat lunch. Which equation is part of the system of equations? A. 3x + 5y = 13 B. 3x + 13 = y C. 9x - 3y = 13 D. 5x + 13 = y

OpenStudy (anonymous):

help!!!

OpenStudy (anonymous):

@bradely